OpenStudy (anonymous):

Find the MLE based on a random sample of size n for the following distribution N (0,theta)

OpenStudy (zarkon):

what have you tried?

OpenStudy (zarkon):

I assume you can find the likelihood function.

OpenStudy (anonymous):

i really dont even know where to start

OpenStudy (zarkon):

do you know the pdf?

OpenStudy (zarkon):

\[f(x;\theta)=\frac{1}{\sqrt{2\pi\theta}}e^{-\frac{x^2}{2\theta}}\]

OpenStudy (zarkon):

now you need to find the likelihood function \[L(\theta)=\prod_{i=1}^{n}\frac{1}{\sqrt{2\pi\theta}}\exp\left(-\frac{x_i^2}{2\theta}\right)\]

OpenStudy (anonymous):

and how do i go about doing that?

OpenStudy (zarkon):

the above is the likelihood function \[L(\theta):=\prod_{i=1}^{n}f(x_i;\theta)\]

OpenStudy (zarkon):

have you done any of these before?

OpenStudy (anonymous):

no im soo lost

OpenStudy (zarkon):

do you understand my notation?

OpenStudy (anonymous):

yea i know what the notations are

OpenStudy (zarkon):

the above is the definition of the likelihood function...you then need to take the log of that and then differentiate with respect to \(\theta\)

OpenStudy (anonymous):

ayi ayi lol

OpenStudy (zarkon):

\[\ell(\theta)=Log(L(\theta))=Log\left(\prod_{i=1}^{n}\frac{1}{\sqrt{2\pi\theta}}\exp\left(-\frac{x_i^2}{2\theta}\right)\right)\]

OpenStudy (anonymous):

what does the exp mean?

OpenStudy (zarkon):

\[\exp(x)=e^{x}\]

OpenStudy (zarkon):

\[=\sum_{i=1}^{n}Log\left(\frac{1}{\sqrt{2\pi\theta}}\exp\left(-\frac{x_i^2}{2\theta}\right)\right)\]

OpenStudy (zarkon):

can you simplify this using the properties of logs...then differentiate. here Log is the natural log

OpenStudy (anonymous):

ok yea i was gunna say isnt it the natural log

OpenStudy (zarkon):

it doesn't have to be...but it make the simplification easier.

OpenStudy (anonymous):

ok.and no im terrible at differentiation

OpenStudy (zarkon):

can you simplify it?

OpenStudy (zarkon):

\[Log(AB)=Log(A)+Log(B)\] \[Log(A^r)=r\cdot Log(A)\]

OpenStudy (anonymous):

i know log (b)

OpenStudy (anonymous):

its just the exponent lol

OpenStudy (anonymous):

right?

OpenStudy (zarkon):

yes..the Log of e is just the exponent

OpenStudy (anonymous):

and log A becomes ?

OpenStudy (zarkon):

\[=\sum_{i=1}^{n}\left[Log\left(\frac{1}{\sqrt{2\pi\theta}}\right)+Log\left(\exp\left(-\frac{x_i^2}{2\theta}\right)\right) \right]\] \[ =\sum_{i=1}^{n}\left[-\frac{1}{2}Log\left(2\pi\theta\right)-\frac{x_i^2}{2\theta} \right]\]

OpenStudy (anonymous):

now i have to differentiate that?

OpenStudy (zarkon):

\[=-\frac{n}{2}Log\left(2\pi\theta\right)-\frac{1}{2\theta}\sum_{i=1}^{n}x_i^2\] then differentiate

OpenStudy (anonymous):

and of course ihave no idea how to do that..im sorry thank you so much more helping me

OpenStudy (zarkon):

treat everything as a constant except \(\theta\)

OpenStudy (zarkon):

after taking the derivative you should get \[-\frac{n}{2\theta}+\frac{1}{2\theta^2}\sum_{i=1}^{n}x_i^2\]

OpenStudy (anonymous):

and is that the final answer?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

what do i have to do next?

OpenStudy (zarkon):

set equal to zero and solve for \(\theta\)

OpenStudy (anonymous):

ok and that gives you the answer?

OpenStudy (zarkon):

it does...though you should really check to see if it is indeed a maximum.

OpenStudy (anonymous):

and what if its not?

OpenStudy (zarkon):

it will be

OpenStudy (anonymous):

i get confused on how to manipulate a function with summations in it

OpenStudy (anonymous):

im sorry you prob think im so dumb..i really just dont understand this

OpenStudy (anonymous):

so theta equals summation xi2/n?

OpenStudy (zarkon):

yes \[\hat{\theta}=\dfrac{\displaystyle\sum_{i=1}^{n}x_i^2}{n}\]

OpenStudy (anonymous):

thank you soo much!!!